How do you work out this problem?

The answer I was given is choice B.1) A sample of 68 books has a mean cost of $ 70.27 and a standard deviation of $ 7.83. Use this information and the special c
Answer #1

Answer)

According to the chebyshev's theorem

At least 1 - (1/k^2) lies with in k standard deviation

1)

51/68 = 0.75 = 75%

1 - (1/2^2) = 75%

So k = 2

Answer is 70.27 - (2*7.83) and 70.27 + (2*7.83)

Between 54.61 and 85.93

2)

46.78 = mean - (3*s.d)

93.76 = mean + (3*s.d)

So k = 3

Answer is 1 - (1/3^2) = 89%

89% of 68 = 61